OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=6/7, (8,-3) The equation of the line in slope-intercept form is y=?

OpenStudy (anonymous):

use the equation y-y1=m(x-x1) Where x1 is 8 and y1 is -3.

OpenStudy (anonymous):

Soo that means the equation would be y=6/7x+4?

OpenStudy (anonymous):

Im confused.. can u explain?

OpenStudy (anonymous):

anyone help ?

OpenStudy (rosey):

u should b a fan of me :) oh and i actually know how to do this haha

OpenStudy (anonymous):

Okay lol.

OpenStudy (rosey):

so the equation would be y - -3 = 6/7 (x - 8)

OpenStudy (rosey):

so just do it from there :)